# A variational formulation of the free energy of mixed quantum-classical systems: coupling classical and electronic density functional theories

### v4

@@ Abstract
Combining classical density functional theory (cDFT) with quantum mechanics (QM) methods offers a computationally efficient alternative to traditional QM/molecular mechanics (MM) approaches for modeling mixed quantum-classical systems at finite temperatures. However, both QM/MM and QM/cDFT rely on somewhat ambiguous approximations.
This paper aims to develop an exact theoretical framework that allows us to clarify the approximations involved in the QM/cDFT formulation and serves as a starting point for further development. We therefore establish a comprehensive density functional theory (DFT) framework for mixed quantum-classical systems within the canonical ensemble. We start with the adiabatic equilibrium density matrix for a system comprising $N_{{\text{qm}}}$ quantum and $N_{{\text{mm}}}$ classical particles and, based on this, we propose a variational formulation of the Helmholtz free energy in terms of the full QM/MM density matrix. Taking advantage of permutational symmetry and thanks to constrained-search methods, we reformulate the computation of the Helmholtz free energy using only the quantum and classical one-body densities, thus generalizing both cDFT and electronic DFT (eDFT) for QM/MM systems.
We then reformulate the functional in order to explicitly make appear the eDFT and cDFT Levy-Lieb functionals, together with a new universal correlation functional for QM/MM systems. A mean-field approximation is finally introduced in the context of solvation problems and we discuss its connection with several existing mixed cDFT-eDFT schemes.
@@ I Introduction
Although electronic structure calculations at zero temperature can routinely handle systems with hundreds to even thousands of atoms, accounting for finite-temperature effects becomes crucial in many applications. Such calculations must therefore be set in a statistical ensemble, where the leading-order entropic contribution arises from nuclear motion. In this context, it is natural to compute the forces acting on the nuclei at the quantum level and then to solve the classical equations of motion to propagate the trajectory. This forms the basis of ab-initio molecular dynamics Iftimie, Minary, and Tuckerman (2005); Silvestrelli and Parrinello (1999). Despite its considerable achievements, it is computationally demanding due to the extensive number of QM calculations required to sample the exponentially large phase space. The convergence rate of the statistics can also vary significantly depending on the targeted properties and conditionsPan et al. (2013); Rozsa et al. (2018). To alleviate ab-initio molecular dynamics computational burden, QM/MM methods were developed, where only a small, critical portion of the system is treated at the QM level, while the remainder is handled at the classical level (MM). The interaction between the classical particles are modeled using parametrized force fields and their dynamics obey Newton’s law of motion. However, even in QM/MM approaches, sampling the phase space of large systems remains computationally intensive.
@@
The efficiency of both eDFT and cDFT stems from a common principle: each method replaces the challenge of computing a high-dimensional object—such as the wavefunction in QM or the probability distribution in MM—with the optimization of a functional that depends only on a low-dimensional quantity, i.e. the one-body density of either quantum electrons or classical molecules. The mathematical framework allowing for such a simplification is the so-called constrained-search principle introduced by Levy and LiebLevy (1979); Lieb (1983) which enables, under certain constraints of representativity, to give an explicit expression of the universal functionals of eDFT and cDFT, although not usable in practical calculations.
@@
Its computational efficiency makes cDFT appealing to model solvent effect at the classical level while retaining a quantum description of the soluteTang et al. (2020); Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024). In this approach, the total Hamiltonian is partitioned into well defined QM and MM Hamiltonians.However, the QM/MM interaction lacks of a unique practical definition. The coupling Hamiltonian is usually described as the sum of an electrostatic interaction combined with a non-bonded term which is often modeled using Lennard-Jones force field to account for repulsion and dispersion effectsSenn and Thiel (2009); Giovannini, Lafiosca, and Cappelli (2017); Clemente, Capece, and Martí (2023). The Mixed cDFT-eDFT approach advantageously replaces the exhaustive phase-space sampling required in QM/MM molecular dynamics with a self-consistent optimization of both classical and quantum functionals. While the idea of merging cDFT and quantum methods is completely natural its theoretical justification is not immediately straightforward.
@@
Petrosyan and co-workersPetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007) justified their so-called joint density-functional formulation by the use of a Levy-Lieb constrained search procedure allowing to express the universal functional as depending on the electron density and the nuclear density rather than the full density matrix describing the electrons and nuclei. To derive a QM/MM formulation, they treat the environment particles as classical by ’integrating out’ the electronic density associated with these nuclei. Howvever, this step makes it difficult to pursue a rigorous derivation and their working functional is rather defined as the sum of the Quantum Kohn-Sham functional describing the electrons of the solute, the classical functional describing the solvent and a coupling functional defined as the difference between the exact functional and the sum of the two former functionals. In other workJeanmairet, Levesque, and Borgis (2020); Tang et al. (2020), a similar form of the functional was invoked without further theoretical justification.
@@
Thus, while QM/cDFT approaches offer potentially attractive computational scaling, their derivation from an exact theory remains somewhat ambiguous. This work addresses this gap by establishing a rigorous mathematical framework that provides a DFT-equivalent description of mixed QM/MM systems. Here, we focus on the canonical ensemble, deriving equations to compute the Helmholtz free energy using only the QM and MM one-body densities. Extending this formulation to the grand canonical ensemble for classical particles is conceptually straightforward but introduces additional complexity in the notations due to the varying number of particles. The main difference between the two ensembles is that, in the grand canonical ensemble, there exists a unique correspondence between the equilibrium density and the external potential. In contrast, in the canonical ensemble, the equilibrium density corresponds to a family of external potentials, defined up to an additive constantLutsko (2022).
Within the present paper, the foundational object from which the QM/MM DFT formulation is derived is the explicit expression of the adiabatic equilibrium density of a mixed QM/MM system in the canonical ensemble. This expression, which was previously established in the adiabatic basis by Nielsen et. al. in the context of QM/MM dynamicsNielsen, Kapral, and Ciccotti (2001), is recovered here following an alternative route (see Sec. II of the supplementary materials). Building on this, we propose a functional formulation of the free energy in terms of general QM and MM high-dimensional densities (see Sec. Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ for a proof of the variational properties). By applying the Levy-Lieb constrained search and assuming additive two-body QM/MM interactions, we then naturally express the free energy in terms of the one-electron density and the classical one-body density. We explicitly establish the connection between eDFT and cDFT through the introduction of a new universal QM/MM correlation one-body density functional. Additionally, we discuss a mean-field approximation and its relationship with previously proposed cDFT/eDFT couplings in the context of QM/MM solvation problems. We also note that this work can be seen as the QM/MM and finite-temperature counterpart of the recently published study by Fromager and LasorneFromager and Lasorne (2024), which formulated an exact DFT framework incorporating both electrons and nuclei at the quantum level—going beyond the Born-Oppenheimer approximation.
The paper is organized as follows. In Sec. II, we begin by discussing the adiabatic equilibrium density for a mixed QM/MM system in the canonical ensemble (Sec. II.1 and II.2). Then, in Sec. II.3, using the explicit form of the equilibrium density, we reformulate the computation of the free energy as a functional minimization over general high-dimensional mixed QM/MM densities. Sec. III focuses on expressing this problem explicitly in terms of the one-body QM and MM densities. In Sec. III.1, we employ the Levy-Lieb constrained search to reformulate the high-dimensional QM/MM functional problem by recognizing a many-to-one correspondence between the high-dimensional QM/MM quantities and the one-body QM and MM densities. In Sec. III.2, we establish a connection between eDFT and cDFT by introducing a universal QM/MM correlation functional of the one-body QM and MM densities. Sec. III.3 then discusses this formulation in the context of solvation problems and recent QM/cDFT coupling approaches. Finally, Sec. IV summarizes the main results and conclusions. The supplementary materials provide an alternative derivation of the equilibrium density for a QM/MM system that does not rely on Liouvillian operators, and the appendix Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ provides a proof of the variational properties of the QM/MM functional.
@@ II Quantum-classical system in the Canonical ensemble: equilibrium density and variational formulation
The main goal of this section is to give a variational formulation for the Helmholtz free energy of a mixed QM/MM system based on a functional minimization over general densities involving $N_{{\text{qm}}}$ and $N_{{\text{mm}}}$ variables. We define in Sec. II.1 our working Hamiltonian describing the QM/MM system. The equilibrium density of a mixed QM/MM system is then given in Sec. II.2, and we briefly discuss the physical meaning of that result. We eventually show in Sec. II.3 that the Helmholtz free energy can be obtained through a variational principle over mixed QM/MM density matrices.
@@ II.1 The Hamiltonian of the quantum-classical system
We adopt the following notation conventions : i) lowercase letters will refer to purely quantum objects (both variables and operators) while capital letters will refer to either purely classical or mixed quantum-classical objects and ii) all bold letters refer to collective variables or operators (i.e. relative to all quantum or classical particles) and non bold letters will refer to individual or pairwise variables or operators.
The total system is divided into a set of $N_{{\text{mm}}}$ identical classical particles, described by their classical positions ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$ and momentum ${\bf P}=\{{{P}}_{1},\ldots,{{P}}_{N_{{\text{mm}}}}\}$, and a set of $N_{{\text{qm}}}$ identical quantum particles described by their position variables ${\bf{q}}=\{{{q}}_{1},\ldots,{{q}}_{N_{{\text{qm}}}}\}$.
The Hamiltonian of the total system is written as
@@
In the rest of the paper, explicit dependency on the QM and MM variables will be dropped for concision when unambiguous. In Eq.(1), $\hat{\bf h}_{{\text{qm}}}$ is the quantum Hamiltonian operator describing $N_{{\text{qm}}}$ quantum identical particles of mass $m$ (typically electrons) in interaction with an external potential $v^{\text{ext}}_{{\text{qm}}}({{q}})$ local in position representation, and interacting among themselves with a pairwise potential ${w}_{{\text{qm}}}({{q}}_{1},{{q}}_{2})$. The explicit form of $\hat{\bf h}_{{\text{qm}}}$ reads then
@@
$\hat{\bf t}=\sum_{i=1}^{N_{{\text{qm}}}}\frac{\hat{p}_{i}^{2}}{2m},\qquad\hat{p}_{i}=-i\hbar\nabla_{{{q}}_{i}},$ (3)
$\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}=\sum_{j=1}^{N_{{\text{qm}}}}v^{\text{ext}}_{{\text{qm}}}({{q}}_{j}),\qquad\hat{\bf w}_{{\text{qm}}}=\sum_{j=1}^{N_{{\text{qm}}}}\sum_{i>j}{w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j}).$ (4)
Then, ${\bf H}_{{\text{mm}}}$ is the classical Hamiltonian function representing $N_{{\text{mm}}}$ identical classical particles of mass $M$, in interaction with a local external potential $V^{\text{ext}}_{{\text{mm}}}({{Q}})$ and interacting among themselves with a local pairwise potential ${W}_{{\text{mm}}}({{Q}}_{i},{{Q}}_{j})$. These particles are typically effective particles such as atoms or rigid molecules instead of the individual electrons composing them and they interact through model, parametrized, force fields. The explicit form of the classical Hamiltonian ${\bf H}_{{\text{mm}}}$ reads then
@@
The function ${W}_{{\text{qm}}}^{{\text{mm}}}({q},{Q})$ represents the effective interaction between an effective particle (such as a nuclei, an atom or a molecule) at position ${Q}$ and a fundamental particle (such as an electron) at position ${q}$.

### v5

@@ Abstract
Combining classical density functional theory (cDFT) with quantum mechanics (QM) methods offers a computationally efficient alternative to traditional QM/molecular mechanics (MM) approaches for modeling mixed quantum-classical systems at finite temperatures. However, both QM/MM and QM/cDFT rely on somewhat ambiguous approximations , the two major ones being: i) the definition of the QM and MM regions as well as the description of their coupling, and ii) the choice of the methods and levels of approximation made to describe each region
This paper addresses the second point and develop an exact theoretical framework that allows us to clarify the approximations involved in the QM/cDFT formulation. We therefore establish a comprehensive density functional theory (DFT) framework for mixed quantum-classical systems within the canonical ensemble. We start by recalling the expression of the adiabatic equilibrium density matrix for a mixed system made of $N_{{\text{qm}}}$ quantum and $N_{{\text{mm}}}$ classical particles. Then, we propose a variational formulation of the Helmholtz free energy in terms of the full , non-equilibrium, QM/MM density matrix. Taking advantage of permutational symmetry and thanks to constrained-search methods, we reformulate the computation of the Helmholtz free energy using only the quantum and classical one-body densities. Therefore, this paper generalizes both cDFT and electronic DFT (eDFT) to QM/MM systems.
We then reformulate the functional to make the standard eDFT and cDFT Levy–Lieb functionals explicitly appear, together with a new universal correlation functional for QM/MM systems. A mean-field approximation is finally introduced in the context of solvation problems and we discuss its connection with several existing mixed cDFT-eDFT schemes. An extension to the semi-grand canonical ensemble, where the number of classical particles is allowed to fluctuate, is provided in the supplementary materials.
@@ I Introduction
Although electronic structure calculations at zero temperature can routinely handle systems with hundreds to even thousands of atoms, accounting for finite-temperature effects becomes crucial in many applications. Such calculations must therefore be set in a statistical ensemble, where the leading-order entropic contribution arises from nuclear motion due to the typically large gaps in electronic energies. In this context, because of the large mass of nuclei with respect to that of electrons, it is physically relevant to maintain electrons at the quantum level and to treat nuclei as classical particles. Nevertheless, the forces acting on the nuclei are computed from the solutions of the quantum electronic ground state, i.e. from the solutions of the Born Oppenheimer Hamiltonian. This forms the basis of ab-initio molecular dynamics Iftimie, Minary, and Tuckerman (2005); Silvestrelli and Parrinello (1999) which recovers the classical nuclear entropic contributions through either classical Monte Carlo algorithms or classical trajectories under the ergotic hypothesis.
Nevertheless, it is important to keep in mind that nuclear quantum effects play a significant role for light nuclei and can substantially alter the rates of certain chemical processes, such as proton transfer reactionsBała et al. (1996); Hanna and Kapral (2005); Kim and Hammes-Schiffer (2006); Costentin, Robert, and Savéant (2010); Hammes-Schiffer and Stuchebrukhov (2010). Moreover, such proton transfers can also be strongly influenced by the solvation effects, specially when polar solvents are usedKapral (2015). In this context, two fundamentally different yet interconnected nuclear quantum effects are at work: (i) the intrinsic delocalization of nuclear wave packets due to the quantum nature of nuclei, and (ii) the nonadiabatic couplings that arise between various adiabatic electronic states. Nuclear delocalization can be captured through explicit nuclear wave functions (see, for example, Ref. Beck et al., 2000 for a review of time-dependent mean-field approaches, and Ref. Hammes-Schiffer, 2021 for the nuclear-electron orbital approach) or via path-integral formalisms such as ring-polymer molecular dynamicsHabershon et al. (2013). Nonadiabatic couplings, on the other hand, are commonly addressed using surface hopping methods based on classical nuclear trajectoriesTully (1990); Hammes-Schiffer and Tully (1994); Ben-Nun and Martı’nez (2000); Li et al. (2005); Richardson and Thoss (2013); Ananth (2013), modified time-dependent mean-field techniquesWorth et al. (2008), semiclassical expansions of the density matrix and its Wigner representationWorth et al. (2008), or even hydrodynamic formulations of quantum mechanics à la Bohm-de BroglieBurghardt and Cederbaum (2001a, b); Burghardt and Bagchi (2006). In contrast, the present work adopts a fully classical treatment of the solvent molecules and focuses on clarifying the formally exact framework for a given QM/MM partitioning, as typically addressed in ab initio molecular dynamics.
Despite its considerable achievements, ab-initio molecular dynamics is computationally demanding due to the extensive number of QM calculations required to sample the exponentially large phase space. The convergence rate of the statistics can also vary significantly depending on the targeted properties and conditionsPan et al. (2013); Rozsa et al. (2018). To alleviate ab-initio molecular dynamics computational burden, QM/MMWarshel and Levitt (1976); Mulholland, Lyne, and Karplus (2000); Senn and Thiel (2009a); Groenhof (2013) methods were developed, where only a small, critical portion of the system is treated at the QM level, while the remainder is handled at the classical level (MM). The interaction between the classical particles are modeled using parametrized force fields and their dynamics obey Newton’s law of motion. The choice of the classical model to represent molecular entities together with the choice of the QM/MM interaction lacks of a unique definition and consists into a whole field of researchRiniker (2018); Harrison et al. (2018); Unke et al. (2021); Wu et al. (2023); Böselt, Thürlemann, and Riniker (2021); Ho et al. (2025) involving both physically-motivated and machine-learning oriented approaches. However, despite the significant gain in computational cost with respect to ab initio molecular dynamics, sampling the phase space of large systems remains computationally intensive even in QM/MM approaches.
@@
The efficiency of both eDFT and cDFT stems from a common principle: each method replaces the challenge of computing a high-dimensional object, such as the wavefunction in QM or the probability distribution in MM, with the optimization of a functional that depends only on a low-dimensional quantity, i.e. the one-body density of either quantum electrons or classical molecules. The mathematical framework allowing for such a simplification is the constrained-search principle introduced by Levy and LiebLevy (1979); Lieb (1983) which enables, under certain constraints of representativity, to give an explicit expression of the universal functionals of eDFT and cDFTDwandaru and Schmidt (2011), although not usable in practical calculations.
@@
Its computational efficiency makes cDFT appealing to model solvent effect at the classical level while retaining a quantum description of the soluteTang et al. (2020); Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024). In this approach, the total Hamiltonian is partitioned into well defined QM and MM Hamiltonians. However, just as in molecular dynamics based QM/MM approaches, the coupling interaction between the QM and MM regions lacks of a unique practical definition. The coupling Hamiltonian is usually described as the sum of an electrostatic interaction combined with a non-bonded term which is often modeled using Lennard-Jones force field to account for repulsion and dispersion effectsSenn and Thiel (2009b); Giovannini, Lafiosca, and Cappelli (2017); Clemente, Capece, and Martí (2023). The Mixed cDFT-eDFT approach advantageously replaces the exhaustive phase-space sampling required in QM/MM molecular dynamics with a self-consistent optimization of both classical and quantum functionals. While the idea of merging cDFT and quantum methods is completely natural its theoretical justification is not immediately straightforward.
@@
Petrosyan and co-workersPetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007) introduced the joint density functional theory which consists in a DFT formulation of the QM/MM problem. The authors obtain a multi component electron/nuclei/solvent DFT formulation, which is justified by the use of a Levy-Lieb constrained search procedure allowing to express the universal functional as depending on the electron density and the nuclear density rather than the full density matrix describing the electrons and nuclei. However, to derive a QM/MM formulation, they treat the environment particles as classical by ’integrating out’ the electronic density associated with these nuclei, without explicitly indicating how this is done. Therefore, this step makes it difficult to pursue a rigorous derivation. Their working functional is instead defined as the sum of the Quantum Kohn-Sham functional describing the electrons of the solute, the classical functional describing the solvent and a coupling functional defined as the difference between the exact functional and the sum of the two former functionals. In other workJeanmairet, Levesque, and Borgis (2020); Tang et al. (2020), a similar form of the functional was invoked without further theoretical justification.
@@
Thus, while QM/cDFT approaches offer potentially attractive computational scaling and appears as a quite natural outcome, their derivation from an exact theory remains somewhat ambiguous. This work addresses this gap by establishing a rigorous mathematical framework that provides a DFT-equivalent description of mixed QM/MM systems. Here, we focus on the canonical ensemble, deriving equations to compute the Helmholtz free energy using only the QM and MM one-body densities. Extending this formulation to the grand canonical ensemble for classical particles is conceptually straightforward but introduces additional complexity in the notations due to the varying number of particles. We therefore do not include it in the core of the paper, but rather in the supplementary materials. This derivation follows a similar route to its formally lighter canonical counterpart, which is reported in the present paper. The main difference between the two ensembles is the existence of a unique mapping between the equilibrium density and the external potential in the grand canonical ensemble. In contrast, in the canonical ensemble, the equilibrium density is associated to a family of external potentials, defined up to an additive constantGonzález et al. (1997); Lutsko (2022).
Our derivation begins with two key assumptions: (a) the entire system is partitioned into a QM region and a MM region; and (b) specific models are available for both the MM subsystem and the QM/MM coupling. As such, we do not address the QM/MM crossover or the accuracy of the MM and QM/MM models themselves. We assume these models are given, as is standard in practical QM/MM calculations.
Under these assumptions, the primary objective of this paper is to derive, in a straightforward manner, the one-body density functional theory (DFT) formulation for a QM/MM system. Therefore, The key steps of this derivation are as follows: i) We begin by defining the general Hamiltonian of the system, considering a broadly applicable QM/MM Hamiltonian (see Sec. II.1). ii) We then express the exact solution of the problem using full-system quantities (i.e., those describing all particles). This includes the Helmholtz free energy of the QM/MM system, along with the corresponding equilibrium density matrix and partition function (see Sec. II.2). iii) Next, we reformulate the problem as the minimization of a functional over generic full-system density matrices. In our case, this means expressing the Helmholtz free energy as a functional of QM/MM density matrices that describe all particles in the system (see Sec. II.3.1). iv) We then decompose this functional into intrinsic (arising from kinetic energy, inter-particle interactions and entropic contributions) and external (arising from external potentials) contributions (see Sec II.3.2). v) Under certain assumptions about the form of the external potentials, we show how the external part of the functional can be written as a linear functional of one-body densities. In our case, this results in a functional of both classical and quantum one-body densities (see Sec. III.1). vi) Leveraging the many-to-one correspondence between full-system and one-body quantities, we perform a Levy–Lieb constrained search to rewrite the intrinsic free energy functional as a universal functional of the one-body densities (see Sec. III.1). vii) We then establish a connection between electronic DFT (eDFT) and classical DFT (cDFT) by introducing a universal QM/MM correlation functional of the one-body QM and MM densities (see Sec. III.2). viii) Finally, we discuss this formulation in the context of solvation problems and recent developments in QM/cDFT coupling approaches (see Sec. III.3). The Appendix Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ contains a proof of the variational properties of the QM/MM functional introduced here, and the Supplementary Material provides a recall of the Von-Neuman variational principle in statistical QM for the canonical ensemble, an alternative derivation of the QM/MM equilibrium density previously established in the adiabatic basis by Nielsen et. al. in the context of QM/MM dynamicsNielsen, Kapral, and Ciccotti (2001), and an extension to the semi grand canonical ensemble of the present formalism. As this work is intended for both the QM and MM communities, which often differ in terminology and conceptual frameworks, we highlight certain elements throughout the text to aid accessibility. Some of these clarifications may appear self-evident to some readers; we hope, however, that they do not unduly burden the readability of the article.
@@ II Quantum-classical system in the Canonical ensemble: equilibrium density and variational formulation
The objective of this section is to give a variational formulation for the Helmholtz free energy of a mixed QM/MM system based on a functional minimization over general densities involving $N_{{\text{qm}}}$ and $N_{{\text{mm}}}$ variables. This is essentially a QM/MM equivalent to the Gibbs variational principle in classical statistical physics or the Von-Neumann variational principle in quantum statistical physics (see the supplementary information for a short review). This variational principle, although hardly usable in practice due to the high-dimensionality of the objects it manipulates, is an important milestone to properly obtain the one-body formulation, which is the aim of this paper.
We define in Sec. II.1 the working Hamiltonian that describes the QM/MM system. Again, we do not discuss the validity of the partitioning or the quality of the coupling term, we assume that these preliminary considerations have already been addressed. The equilibrium density of a mixed QM/MM system is then given in Sec. II.2, and we briefly discuss the physical meaning of that result. We eventually show in Sec. II.3 that the Helmholtz free energy can be obtained through a variational principle over mixed QM/MM density matrices.
@@ II.1 The Hamiltonian of the quantum-classical system
We consider a QM/MM system, where the partition and the choice of the model to describe the MM part and the coupling term have already been made
We adopt the following notation conventions: (i) lowercase letters denote purely quantum quantities (variables or operators), while uppercase letters denote classical or mixed quantum–classical quantities; (ii) bold symbols denote collective quantities (over all particles), while non-bold symbols denote single-particle or pairwise quantities. The only exception is the number of quantum particles, labeled $N_{{\text{qm}}}$. Atomic units are used throughout.
The total system is divided into a set of $N_{{\text{mm}}}$ identical classical particles, described by their classical positions ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$ and linear momentum ${\bf P}=\{{{P}}_{1},\ldots,{{P}}_{N_{{\text{mm}}}}\}$ and a set of $N_{{\text{qm}}}$ identical quantum particles described by their position variables ${\bf{q}}=\{{{q}}_{1},\ldots,{{q}}_{N_{{\text{qm}}}}\}$.
The total Hamiltonian is written as the sum of the QM and MM Hamiltonians, and a QM/MM coupling term, namely
@@
In the rest of the paper, explicit dependency on the QM and MM variables will be dropped for concision when unambiguous. In Eq.(1), $\hat{\bf h}_{{\text{qm}}}$ is the quantum Hamiltonian operator describing $N_{{\text{qm}}}$ quantum identical particles of mass $m$ (typically electrons). These particles interact with an external potential $v^{\text{ext}}_{{\text{qm}}}({{q}})$, which is local in the position representation, and with each other through a pairwise potential ${w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j})$. The explicit form of $\hat{\bf h}_{{\text{qm}}}$ reads then
@@
$\hat{\bf t}=\sum_{i=1}^{N_{{\text{qm}}}}\frac{\hat{p}_{i}^{2}}{2m},\qquad\hat{p}_{i}=-i\nabla_{{{q}}_{i}},$ (3)
$\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}=\sum_{i=1}^{N_{{\text{qm}}}}v^{\text{ext}}_{{\text{qm}}}({{q}}_{i}),\qquad\hat{\bf w}_{{\text{qm}}}=\sum_{j=1}^{N_{{\text{qm}}}}\sum_{i>j}{w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j}).$ (4)
Then, ${\bf H}_{{\text{mm}}}$ is the classical model Hamiltonian function representing $N_{{\text{mm}}}$ identical particles of mass $M$, in interaction with a local external potential $V^{\text{ext}}_{{\text{mm}}}({{Q}})$ and interacting among themselves through a local pairwise potential ${W}_{{\text{mm}}}({{Q}}_{i},{{Q}}_{j})$. These are typically effective particles, such as atoms or rigid molecules, and they interact with each other through parametrized force fields. The explicit form of the classical Hamiltonian ${\bf H}_{{\text{mm}}}$ reads then
@@
The function ${W}_{{\text{qm}}}^{{\text{mm}}}({q},{Q})$ represents the effective interaction between an effective particle (such as a nuclei, an atom or a molecule) at position Q{Q} and a fundamental particle (such as an electron) at position q{q}.